# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v6

@@ 4.1.2 Implementation Details
Following the preprocessing method used in MEGraph [26], the 68 facial landmarks undergo a conversion into 49 facial internal landmarks, excluding those associated with the facial contour. Subsequently, the images undergo augmentation through similarity transformation, encompassing in-plane rotation, uniform scaling, and translation, utilizing the 49 facial landmarks. The resulting images are standardized to a resolution of 256$\times$256$\times$3 pixels and then randomly cropped into 224$\times$224$\times$3 pixels, with the addition of a random horizontal flip.
For the pre-training process, the encoder is conducted on one NVIDIA GeForce RTX 3090 GPU with 24GB for one day and implemented by PyTorch using an AdamW optimizer with $\beta_{1}$= 0.9, $\beta_{2}$= 0.999 and weight decay of $10^{-6}$. We set a learning rate of 10-5 and 60 epochs. In our discriminative contrastive learning framework, where sample pairs are selected within the same batch, the batch size significantly impacts the experimental results. For consistency, we have set the batch size to 32 in our experiments. In terms of the hyperparameters in AUNCE, $n_{au}$ is set as 12 for BP4D, BP4D+, Aff-Wild2, 10 for GFT, and 8 for DISFA respectively, and the temperature coefficient $\tau$ is 0.5. The analysis for $\beta$ and $p_{1}$, $p_{2}$, $p_{3}$ will introduce in Section. 4.4.1.
For the linear evaluation protocol [9, 13], we extract $n_{au}$ 2048-dimensional feature vectors from the model pre-trained using the AUNCE loss. A linear classifier is then trained on these features using AdamW optimizer with the same settings as the pre-training phase, a batch size of 256 on a single GPU, a learning rate of 10-3, and 100 epochs. The results are obtained through training with the weighted cross-entropy loss (WCE), and the highest test accuracy across our models serves as a surrogate measure for the quality of the learned representations.
@@ 4.1.3 Evaluation Metrics
Aligned with current state-of-the-art methodologies, we evaluate our AU detection method using three frame-based metrics: F1-macro score, F1-micro score, and Accuracy. The F1-macro score is obtained by calculating the F1 score for each AU class individually and then averaging these values, which is a standard metric employed by many existing methods. Inspired by the work of Hinduja et al. [33], we also incorporate the F1-micro score, which is computed by aggregating the true positives, false positives, and false negatives across all classes before calculating the overall F1-score. This metric is particularly useful for evaluating performance in imbalanced datasets. Additionally, we include Accuracy, which is defined as the ratio of correctly classified frames to the total number of frames, to provide a more comprehensive assessment of model performance. For simplicity, we will refer to the F1-macro score as F1-score in the remainder of the text.
Table 2: Quantitative comparison results on BP4D dataset, where the evaluation metric is F1-score(%).
@@
Regional feature learning-based methods
EAC-Net [4] 2018 TPAMI 39.0 35.2 48.6 76.1 72.9 81.9 86.2 58.8 37.5 59.1 35.9 35.8 55.9
ROI [3] 2017 CVPR 36.2 31.6 43.4 77.1 73.7 85.0 87.0 62.6 45.7 58.0 38.3 37.4 56.4
ARL [5] 2019 TAC 45.8 39.8 55.1 75.7 77.2 82.3 86.6 58.8 47.6 62.1 47.4 55.4 61.1
AC2[34] 2024 IJCV 54.2 54.7 56.5 77.0 76.2 84.0 89.0 63.6 54.8 63.6 46.5 54.8 64.6
@@
($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$
0.15) 0.12) 0.17) 0.21) 0.19) 0.22) 0.22) 0.17) 0.14) 0.15) 0.14) 0.15) 0.17)
Table 3: Quantitative comparison results on DISFA dataset, where the evaluation metric is F1-score(%).
@@
Regional feature learning-based methods
EAC-Net [4] 2018 TPAMI 41.5 26.4 66.4 50.7 80.5 89.3 88.9 15.6 48.5
ROI [3] 2017 CVPR 41.5 26.4 66.4 50.7 80.5 89.3 88.9 15.6 48.5
ARL [5] 2019 TAC 43.9 42.1 63.6 41.8 40.0 76.2 95.2 66.8 58.7
AC2[34] 2024 IJCV 57.8 59.2 70.1 50.1 54.4 75.1 90.3 66.2 65.4
@@
MMA-Net [37] 2023 PRL 63.8 54.8 73.6 39.2 61.5 73.1 92.3 70.5 66.0
@@
($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$
0.37) 0.22) 0.42) 0.31) 0.36) 0.49) 0.42) 0.28) 0.36)
Table 4: Quantitative comparison results on BP4D+ dataset, where the evaluation metric is F1-score(%).
@@
AC2[34] 2024 IJCV 42.3 35.4 26.7 80.7 87.0 90.9 85.8 73.3 45.3 43.4 50.3 29.0 57.5
AUNCE(Ours) - 42.5 41.5 27.9 84.1 86.5 91.7 91.3 72.1 49.7 48.8 48.0 33.5 59.8
($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$
0.18) 0.46) 0.14) 0.32) 0.23) 0.17) 0.14) 0.49) 0.33) 0.25) 0.12) 0.24) 0.26)
Table 5: Quantitative comparison results on GFT dataset, where the evaluation metric is F1-score(%).
@@
($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$
0.33) 0.18) 0.28) 0.24) 0.41) 0.11) 0.35) 0.23) 0.42) 0.35) 0.29)
Table 6: Quantitative comparison results on Aff-Wild2 dataset, where the evaluation metric is F1-score(%).
@@
($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$ ($\pm$
0.25) 0.37) 0.28) 0.33) 0.12) 0.47) 0.31) 0.17) 0.08) 0.42) 0.39) 0.32) 0.29)
Table 7: Quantitative comparison results for common 10 AUs of cross evaluation between BP4D and GFT. BP4D → GFT denotes training on BP4D and testing on GFT. The evaluation metric is F1-score(%).
Dataset Method AU Index Avg.
@@
BP4D → GFT AC2[34] 28.0 35.7 22.7 70.5 69.2 65.2 16.3 29.0 40.8 45.2 42.3
Ours (WCE) 30.2 37.9 24.8 66.4 72.3 62.6 20.1 28.6 44.6 38.8 42.6
Ours (AUNCE) 32.4 50.1 10.5 71.0 77.6 70.2 22.3 26.5 50.9 40.4 45.2
GFT → BP4D Ours (WCE) 49.5 45.5 26.0 20.4 52.3 36.6 17.3 18.6 50.5 54.4 37.1
AC2[34] 51.9 49.3 25.8 24.6 50.1 40.9 15.3 20.1 47.2 58.3 38.4
Ours (AUNCE) 52.3 50.8 24.6 18.7 58.8 38.8 20.6 22.5 56.6 62.9 40.7
@@ 4.2.1 Evaluation on BP4D
Since we use AU labels to assist the process of selecting positive sample pairs, we introduce five groups of supervised AU detection frameworks to compare with our method. The quantitative results of state-of-the-art methods from the past seven years for BP4D dataset are shown in Table 2. Compared to the latest state-of-the-art results, our method AUNCE demonstrates significantly superior performance.
Methods in the first group are based on feature extraction of regional area around AUs, including EAC-Net [4], ROI [3], ARL [5] and AC2[34]. Our approach achieves a notable improvement, with an increase in F1-score ranging from 1.5% to 10.2% over these methods. This advancement is primarily due to the feature extraction methods of these frameworks neglecting critical correlation information among AUs and failing to specifically address the class imbalance issue for each AU. Furthermore, the precision of the encoders utilized by these methods is generally lower.
@@
Methods in the third group mainly utilize the semantic prior knowledge, including MMA-Net [37] and SEV-Net [38]. Our approach demonstrates an average improvement of 2.7% and 2.2%, respectively. This phenomenon can be attributed to the inherent limitations of these methods in effectively capturing and representing the distinct information encoded within AUs using simple semantic descriptions.
@@ 4.2.2 Evaluation on DISFA
Comparable outcomes are evident in the DISFA dataset, as outlined in Table 3, where AUNCE attains the highest overall performance in AU detection. Furthermore, a comparison between Table 2 and Table 3 reveals that recent methods such as EmoCo, CLP, and MEGraph demonstrate strong performance on the BP4D dataset but achieve only mediocre results on the DISFA dataset. This disparity suggests their limited generalizability. The more pronounced class imbalance issue in DISFA, as shown in Table 1, likely contributes to this performance gap. In contrast, AUNCE consistently delivers robust performance on both BP4D and DISFA datasets and achieves more stable results across various AUs due to our proposed negative sample re-weighting strategy.
@@ 4.2.6 Cross dataset validation of AUNCE
To evaluate the generalization and cross-dataset reliability of our framework, we compare AUNCE with two baselines: one using the same encoder with weighted cross-entropy loss (WCE) and another, AC2[34]. Results are presented in Table 7. Since the 10 AUs in the GFT dataset overlap with the 12 AUs in BP4D, we perform cross-dataset evaluations between BP4D and GFT. For BP4D-to-GFT, we use the three BP4D models trained via 3-fold cross-validation and report the average results. For GFT-to-BP4D, the trained GFT model is tested on the three BP4D test sets, and average results are reported.
As shown in Tables 2 and 5, the performance of AUNCE is notably lower in this cross-dataset evaluation. This performance drop can be attributed to the substantial domain gap between BP4D and GFT. Additionally, our findings show that AUNCE demonstrates superior performance compared to the other two approaches that rely on WCE loss in both BP4D → GFT and GFT → BP4D tasks. This result further underscores the robustness and generalization capability of our proposed discriminative contrastive learning framework.
@@
Backbone Source Loss Dataset
@@ 4.3.1 Compared with weighted cross-entropy loss and weighted focal loss
The primary innovation of this paper is the introduction of AUNCE, a novel contrastive loss designed to be compatible with various AU detection encoders. To assess its effectiveness, we compare the proposed AUNCE loss with traditional weighted cross-entropy loss (WCE) and the recently introduced weighted focal loss (WFL) [7] for AU detection during the pretraining phase, across multiple AU detection encoders. The results, presented in Table 8, demonstrate that AUNCE consistently outperforms the other losses in AU detection. Notably, the MEGraph encoder achieves the highest performance, with scores of 66.1% and 66.4% on BP4D and DISFA datasets, respectively, when trained using AUNCE. This highlights the advantage of emphasizing the capture of distinct features from individual samples, rather than solely relying on pixel-level information of the entire face in relation to AU detection.
Table 9: Comparison of the computational complexity and efficiency of the weighted cross-entropy loss (WCE), the weighted focal loss (WFL), and the proposed AUNCE loss on BP4D and DISFA dataset, where encoders of these three losses are both MEGraph with the first-stage training network.
@@ 4.3.2 Evaluating other components
Weight coefficient $w_{i}$: The weights $w_{i}$ have been pointed out as a fixed value in Section. 3.1, which achieve an increase of 0.9% in F1-macro score, 0.2% in F1-micro score and 0.2% in Accuracy. For an AU, its activation indicator is a binary variable, and the corresponding information content can be measured using entropy: $-\log(p_{i})$. Therefore, AUs with lower activation probabilities carry more information and should be assigned higher weights. This forms the theoretical basis for our weighting strategy.
@@
Negative sample re-weighting strategy: The negative sample re-weighting strategy, which prioritizes minority AUs during training, improves the F1-macro score by 3.1%, the F1-micro score by 0.5%, and accuracy by 0.6% compared to model B, demonstrating its ability to adjust backpropagation gradients for minority and majority classes, enabling faster updates with fewer samples. Additionally, model E achieves a 2.3% average increase in F1-macro score compared to model C, further validating the strategy’s effectiveness. As shown in Fig. 4, model E (top row) focuses more on facial regions relevant to rare AUs, enhancing the detection of these challenging cases and reinforcing its utility in addressing class imbalance.
Figure 5: T-SNE visualization of feature representations on BP4D and DISFA datasets. Colors indicates whether AU1 exists. Top row: (a) Representations optimized by WCE on BP4D dataset, (b)-(f) Visualizations of ablation study on BP4D dataset, respectively represent model A-E. Bottom row: (g) Representations optimized by AUNCE on DISFA dataset, (h)-(l) Visualizations of ablation study on DISFA dataset, respectively represent model A-E.
@@ 4.3.3 $p_{1},p_{2},p_{3},p_{4}$, the probabilities of various positive sample pairs
The experimental results are presented in Table 11. When $p_{1}$=0.15, $p_{2}$=0.15, $p_{3}$=0.7, and $p_{4}$=0, the experimental result is optimal. It is observed that images with the lowest similarity to original images contribute minimally to the training process, as these images essentially represent noise and errors. The mixture of all positive samples plays a pivotal role, as it encourages the feature representation of each instance to be closer to the class centroid, facilitating the capture of essential characteristics of the positive class. The indispensability of the remaining two positive samples lies in their ability to mitigate the impact of noisy and false samples, thereby reinforcing the robustness of the training process.

### v7

@@ 4.1.2 Implementation Details
Following the preprocessing method used in MEGraph [22], the 68 facial landmarks undergo a conversion into 49 facial internal landmarks, excluding those associated with the facial contour. Subsequently, the images undergo augmentation through similarity transformation, encompassing in-plane rotation, uniform scaling, and translation, utilizing the 49 facial landmarks. The resulting images are standardized to a resolution of $\times256$$\times$3 pixels and then randomly cropped into $\times224$$\times$3 pixels, with the addition of a random horizontal flip.
For the pre-training process, the encoder is conducted on one NVIDIA GeForce RTX 3090 GPU with 24GB for one day and implemented by PyTorch using an AdamW optimizer with $\beta_{1}$= 0.9, $\beta_{2}$= 0.999 and weight decay of $10^{-6}$. We set a learning rate of 10-5 and 60 epochs. Our random seed value is set to 3407, consistent with the linear evaluation protocol. In our discriminative contrastive learning framework, where sample pairs are selected within the same batch, the batch size significantly impacts the experimental results. For consistency, we have set the batch size to 32 in our experiments. In terms of the hyperparameters in AUNCE, $n_{au}$ is set as 12 for BP4D, BP4D+, Aff-Wild2, 10 for GFT, and 8 for DISFA respectively, and the temperature coefficient $\tau$ is 0.5. The analysis for $\beta$ and $p_{1}$, $p_{2}$, $p_{3}$ will introduce in Section. 4.4.2. Additionally, data partitioning is performed using 3-fold cross-validation, where the partitioning is subject-independent to avoid identity leakage.
For the linear evaluation protocol [7, 11], we extract $n_{au}$ 2048-dimensional feature vectors from a model pre-trained with the AUNCE loss. A linear classifier is trained using the AdamW optimizer with the same settings as the pre-training phase, including a batch size of 256, a learning rate of 10-3, and 100 epochs. Results are obtained with weighted cross-entropy loss (WCE).
Although formal significance tests (paired t-test, Wilcoxon) are important, most baselines report only average performance without multiple independent runs. To address this, we report standard deviations (computed over at least five runs) in Tables 3–7, showing the robustness of our framework. Our supplementary analysis indicates that in most experimental scenarios, our method’s performance improvement over baselines exceeds twice the standard deviation.
@@ 4.1.3 Evaluation Metrics
We evaluate our AU detection method using three metrics: F1-macro, F1-micro, and Accuracy. The F1-macro score averages the F1 scores for each AU class. The F1-micro score, as proposed by Hinduja et al. [29], aggregates true positives, false positives, and false negatives across all classes, making it suitable for imbalanced datasets. Accuracy is defined as the ratio of correctly classified frames to total frames. We use F1-score to refer to the F1-macro score.
Table 2: Quantitative comparison results on BP4D dataset, where the evaluation metric is F1-score(%). The values following the $\pm$’ symbol denote the standard deviation (SD) of the data.
@@
ETDF [34] 2025 TAC 54.7 50.8 57.1 78.8 79.6 84.6 88.0 67.0 54.9 62.9 48.6 54.5 65.1
Attention mechanism-based methods
ARL [35] 2022 TAC 45.8 39.8 55.1 75.7 77.2 82.3 86.6 58.8 47.6 62.1 47.4 55.4 61.1
AC2[3] 2025 IJCV 54.2 54.7 56.5 77.0 76.2 84.0 89.0 63.6 54.8 63.6 46.5 54.8 64.6
ESCDA [5] 2026 PR 56.5 44.5 59.6 79.4 77.4 84.7 88.7 64.9 57.0 65.0 52.5 55.6 65.5
@@
($\pm$0.15)
($\pm$0.12)
($\pm$0.17)
($\pm$0.21)
($\pm$0.19)
($\pm$0.22)
($\pm$0.22)
($\pm$0.17)
($\pm$0.14)
($\pm$0.15)
($\pm$0.14)
($\pm$0.15)
($\pm$0.17)
Table 3: Quantitative comparison results on DISFA dataset, where the evaluation metric is F1-score(%). The values following the $\pm$’ symbol denote the standard deviation (SD) of the data.
@@
ETDF [34] 2025 TAC 62.6 54.7 70.8 46.3 51.7 76.3 94.4 59.8 64.6
GLEE-Net [37] 2024 TCSVT 61.9 54.0 75.8 45.9 55.7 77.6 92.9 60.0 65.5
Attention mechanism-based methods
ROI [38] 2017 CVPR 41.5 26.4 66.4 50.7 80.5 89.3 88.9 15.6 48.5
ARL [35] 2022 TAC 43.9 42.1 63.6 41.8 40.0 76.2 95.2 66.8 58.7
AC2[3] 2025 IJCV 57.8 59.2 70.1 50.1 54.4 75.1 90.3 66.2 65.4
@@
($\pm 0.37$) ($\pm 0.22$) ($\pm 0.42$) ($\pm 0.31$) ($\pm 0.36$) ($\pm 0.49$) ($\pm 0.42$) ($\pm 0.28$) ($\pm 0.36$)
Table 4: Quantitative comparison results on BP4D+ dataset in terms of cross-dataset evaluation, where the evaluation metric is F1-score(%). The values following the $\pm$’ symbol denote the standard deviation (SD) of the data.
@@
AC2[3] 2025 IJCV 42.3 35.4 26.7 80.7 87.0 90.9 85.8 73.3 45.3 43.4 50.3 29.0 57.5
GLEE-Net [37] 2024 TCSVT 39.8 37.9 41.6 83.4 88.2 90.2 87.4 76.6 48.3 42.9 47.7 29.8 59.5
AUNCE(Ours) - 45.5 44.5 30.9 87.1 89.5 94.7 94.3 72.1 52.7 51.8 52.0 35.5 62.8
($\pm$ 0.25)
($\pm$ 0.37)
($\pm$ 0.28)
($\pm$ 0.33)
($\pm$ 0.12)
($\pm$ 0.47)
($\pm$ 0.31)
($\pm$ 0.17)
($\pm$ 0.08)
($\pm$ 0.42)
($\pm$ 0.39)
($\pm$ 0.32)
($\pm$ 0.29)
Table 5: Quantitative comparison results on GFT dataset, where the evaluation metric is F1-score(%). The values following the $\pm$’ symbol denote the standard deviation (SD) of the data.
@@
($\pm 0.33$) ($\pm 0.18$) ($\pm 0.28$) ($\pm 0.24$) ($\pm 0.41$) ($\pm 0.11$) ($\pm 0.35$) ($\pm 0.23$) ($\pm 0.42$) ($\pm 0.35$) ($\pm 0.29$)
Table 6: Quantitative comparison results on Aff-Wild2 dataset, where the evaluation metric is F1-score(%). The values following the $\pm$’ symbol denote the standard deviation (SD) of the data.
@@
($\pm$ 0.25)
($\pm$ 0.37)
($\pm$ 0.28)
($\pm$ 0.33)
($\pm$ 0.12)
($\pm$ 0.47)
($\pm$ 0.31)
($\pm$ 0.17)
($\pm$ 0.08)
($\pm$ 0.42)
($\pm$ 0.39)
($\pm$ 0.32)
($\pm$ 0.29)
[…]
Dataset Method Source AU Index Avg.
@@
[…]
AC2[3] 2025 IJCV 28.0 35.7 22.7 70.5 69.2 65.2 16.3 29.0 40.8 45.2 42.3
Ours (WCE) - 30.2 37.9 24.8 66.4 72.3 62.6 20.1 28.6 44.6 38.8 42.6
Ours (AUNCE) - 32.4 50.1 10.5 71.0 77.6 70.2 22.3 26.5 50.9 40.4 45.2
[…]
AC2[3] 2025 IJCV 51.9 49.3 25.8 24.6 50.1 40.9 15.3 20.1 47.2 58.3 38.4
D2CA [20] 2025 TIP 31.0 33.0 43.2 60.5 53.7 71.1 21.1 36.5 36.4 44.6 43.1
Ours (AUNCE) - 55.3 53.8 25.4 23.9 61.8 38.8 23.6 25.5 59.6 65.9 43.7
@@ 4.2.1 Evaluation on BP4D
Since we use AU labels to assist the process of selecting positive sample pairs, we compare our approach with five groups of supervised AU detection frameworks. For the first four categories, we select three recent or representative methods each, while the final category, contrastive learning-based approaches, is examined in greater detail. The quantitative results of state-of-the-art methods from the past five years for BP4D dataset are shown in Table 2. Compared to the latest state-of-the-art results, our method AUNCE demonstrates significantly superior performance.
@@
Methods in the second group mainly utilize the semantic prior knowledge, including MMA-Net [32], SEV-Net [33] and ETDF [34]. Our approach demonstrates an average improvement of 2.7%, 2.2% and 1.0%, respectively. This phenomenon can be attributed to the inherent limitations of these methods in capturing the fine-grained and localized cues associated with individual AUs, as simple semantic descriptions often fail to reflect the underlying visual complexity. Additionally, their dependence on predefined semantic priors may introduce inductive biases that hinder the model’s ability to flexibly adapt to the nuanced and heterogeneous nature of facial behaviors across individuals and contexts.
Methods in the third group leverage attention mechanisms to selectively focus on regional facial areas surrounding AUs, including ARL [35], AC2[3] and ESCDA [5]. Our approach achieves a notable improvement, with an increase in F1-score ranging from 0.6% to 5.0% over these methods. This advancement is primarily due to the feature extraction methods of these frameworks neglecting critical correlation information among AUs and failing to specifically address the class imbalance issue for each AU. Furthermore, their reliance on attention mechanisms may lead to suboptimal encoding when attention is misdirected or insufficiently localized, limiting the precision of AU-specific feature representations.
@@ 4.2.2 Evaluation on DISFA
Comparable results are observed on the DISFA dataset (Table 3), where AUNCE achieves the highest overall AU detection performance, though the improvement margin is smaller than on BP4D (Table 2). This discrepancy can be attributed to the more severe class imbalance, annotation noise, and uneven AU distribution in DISFA, which collectively hinder model generalization. Nevertheless, unlike methods such as MEGraph [22], CLP [9], and SEV-Net [33] that perform well on BP4D but degrade noticeably on DISFA, AUNCE maintains stable and superior performance across both datasets, benefiting from the proposed negative sample re-weighting strategy that mitigates the effects of noisy and imbalanced samples and enhances robustness across diverse AUs.
@@ 4.2.6 Cross dataset validation of AUNCE
To assess the generalization capability and cross-dataset robustness of our framework, we compare AUNCE against three baselines: a model using the same encoder with weighted cross-entropy loss (WCE), AC2[3], and D2CA [20], the latter of which is specifically designed for cross-domain AU detection. The comparative results are summarized in Table 7. Given that 10 AUs in the GFT dataset overlap with 12 AUs in BP4D, we conduct cross-dataset evaluations between BP4D and GFT.
For the BP4D-to-GFT setting, we utilize three BP4D models trained via 3-fold cross-validation and report their averaged performance. Conversely, for the GFT-to-BP4D setting, the model trained on GFT is evaluated on the three BP4D test folds, with average results reported accordingly.
As shown in Tables 2 and 5, AUNCE exhibits a performance drop in cross-dataset settings, largely due to the significant domain shift between BP4D and GFT. Nevertheless, AUNCE consistently outperforms all three baselines in both BP4D-to-GFT and GFT-to-BP4D scenarios, underscoring its superior generalization ability and robustness across domains.
@@
Encoder Source Loss Dataset
@@ 4.3.1 Compared with weighted cross-entropy loss and weighted focal loss
The primary innovation of this paper is the introduction of AUNCE, a novel contrastive loss designed to be compatible with various AU detection encoders. To evaluate its effectiveness, we compare the proposed AUNCE loss against the traditional weighted cross-entropy loss (WCE) and the recently introduced weighted focal loss (WFL) [10] during the pretraining stage, using multiple AU detection encoders. While numerous loss functions have been proposed to address sample imbalance, WFL has emerged as a widely adopted and empirically validated approach for mitigating positive–negative imbalance in AU detection. The results, presented in Table 8, demonstrate that AUNCE consistently outperforms the other losses in AU detection. Notably, the MEGraph encoder achieves the highest performance, with scores of 66.1% and 66.4% on BP4D and DISFA datasets, respectively, when trained using AUNCE. This highlights the advantage of emphasizing the capture of distinct features from individual samples, rather than solely relying on pixel-level information of the entire face in relation to AU detection.
Table 9: Comparison of the computational complexity and efficiency of the weighted cross-entropy loss (WCE), the weighted focal loss (WFL), and the proposed AUNCE loss on BP4D and DISFA dataset, where encoders of these three losses are both MEGraph with the first-stage training network. Labels denote ”N = samples” and ”C = classes”.
@@ 4.3.2 Evaluating other components
Weight coefficient $w_{i}$: The weights $w_{i}$ have been pointed out as a fixed value in Section. 3.1, which achieve an increase of 0.9% in F1-macro score, 0.2% in F1-micro score and 0.2% in Accuracy. For an AU, its activation indicator is a binary variable, and the corresponding information content can be measured using entropy: $\log(p_{i})$. Therefore, AUs with lower activation probabilities carry more information and should be assigned higher weights. This forms the theoretical basis for our weighting strategy.
@@
Negative sample re-weighting strategy: The negative sample re-weighting strategy, which emphasizes minority AUs during training, improves the F1-macro score by 3.1%, the F1-micro score by 0.5%, and overall accuracy by 0.6% compared to model B. This demonstrates its effectiveness in adjusting backpropagation gradients between majority and minority classes, allowing more efficient learning from limited samples. Furthermore, model E shows a 2.3% average improvement in F1-macro score over model C, with notable gains in several highly imbalanced AUs, including AU1, AU2, AU4, AU15, AU17, AU23, and AU24, further validating the strategy’s impact. To qualitatively assess its effectiveness, we additionally provide visualization results in Fig. 4. As shown, model E (top row) attends more closely to facial regions associated with rare AUs, enhancing detection performance in these difficult cases and reinforcing the method’s capacity to mitigate class imbalance.
Figure 5: T-SNE visualization of feature representations on BP4D and DISFA datasets. Colors indicates whether AU1 exists. Green (#5BAB54) points denote negative samples (Label = 0), and purple (#885F9C) points denote positive samples (Label = 1). Top row: (a) Representations optimized by WCE on BP4D dataset, (b)-(f) Visualizations of ablation study on BP4D dataset, respectively represent model A-E. Bottom row: (g) Representations optimized by AUNCE on DISFA dataset, (h)-(l) Visualizations of ablation study on DISFA dataset, respectively represent model A-E.